OpenStudy (anonymous):

Chuck's CheeseHouse makes two types of cheese; mozzarella and parmesan. Mozzarella costs $2.60 per pound and parmesan costs $3.30 per pound. How much parmesan cheese went into making a 60 pound mixture worth $3.10 per pound? Round to the nearest pound. Answer 17 lbs 23 lbs 43 lbs 37 lbs

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@mickey4691

OpenStudy (anonymous):

@ktnguyen1

OpenStudy (anonymous):

@ktnguyen1 HELP ME OUT!

OpenStudy (anonymous):

@ktnguyen1 well man you tried

OpenStudy (anonymous):

Im still thinking. this make me confused too

OpenStudy (anonymous):

i got 43

OpenStudy (phi):

These mixture problems can be solved using 2 equations. First, pick two variable names for the two different ingredients x for pounds of mozzarella y for pounds parmesan Second, look for an equation telling us how these two add up. Read: How much parmesan cheese went into making a 60 pound mixture worth $3.10 per pound? 60 pound mixture is the clue. Both cheeses together add to 60 lbs. write this as an equation: x+y= 60 Third, look for an equation relating the cost. They tell us the cost per pound of both cheeses. The cost of the mozzarella is 2.60 * x (cost per pound times number of pounds) The cost of the parmesan is 3.30 * y The cost of the mixture is 60 lbs * 3.10 Now use the idea that the cost of the mozzarella plus the cost of the parmesan adds up to the cost of the mixture. As an equation this is 2.60x + 3.30y= 60*3.10 You now have 2 equations and 2 unknowns x+y= 60 2.60x + 3.30y= 186 Can you solve these?